# Dynamical Lensing Tomography of Black Hole Ringdown

### v2

@@ III Lensing Tomography
Figure 3: Deflection angle of geodesics terminating at infinity, defined as the grand arc difference with the unperturbed direction, plotted as a function of the observation time $t_{0}-r_{*}(r_{o})$. Different colors represent geodesics launched with various impact parameter shifts $\Delta\tilde{\rho}\equiv\tilde{\rho}-\tilde{\rho}_{c}$ from the critical curve in Kerr spacetime, where $\tilde{\rho}_{c}\approx 5.04M$, along the $+\tilde{X}$-axis. The BH, observer, and ringdown configuration are consistent with the setup described in Fig. 1. Near the critical curve, we find an oscillatory, exponentially decaying dependence at intermediate times, in very good (to sub-percent level) agreement with the ringdown of the GW itself. At later times, a pattern of slow decay is observed, consistent with a power law $\tilde{\delta}\sim(t_{o}-r_{*})^{-p}$ with $p\sim 3$.
On the other hand, for late behavior and cases with large impact parameters $\Delta\tilde{\rho}/M=5,10,15$, the variations exhibit a slow decay without oscillations. The two different types of behavior is attributable to the deflections developing in two distinct regions. From the BH to the observer in the $+t$ direction, the GW and photons co-propagate, so the deflection that develops along this path reflects the ringdown metric perturbations present when the photon starts to co-propagate with the GWs. Before entering the co-propagation region, particularly as the photon travels from infinity towards the BH, the deflection received along the path is predominantly influenced by the time it passes through the filter at $t\approx r_{*}$. For smaller impact parameters, the co-propagation duration lasts longer, resulting in deflected angles starting with an exponential decay and oscillation, mirroring the GW ringdown. However, in late-time scenarios and with larger impact parameters, the deflection is primarily influenced by non-co-propagating impacts after $t\approx r_{*}$. The case with $\Delta\tilde{\rho}=15M$ shows larger deflection than those with $\Delta\tilde{\rho}=5M$ and $10M$ because its polar angle $\theta$ at $t\approx r_{*}$ is closer to $\pi$, where the metric perturbation for the $\ell=m=2$ mode is most dominant. Our results are consistent with a power law, $\tilde{\delta}\sim(t_{o}-r_{*})^{-p}$, where $p\approx 3$. This scaling behavior resembles that observed in GW lensing from localized sources Damour and Esposito-Farese (1998); Kopeikin et al. (1999); Kopeikin and Schaefer (1999); Kopeikin et al. (2006), where the dominant contribution to lensing at distances far from the source is proportional to $h_{\mu\nu}p^{\mu}p^{\nu}$, which is only non-vanishing for the $\mathcal{O}(1/r^{3})$ component of the metric perturbations.
Figure 4: Similar to Fig. 3, this figure displays variations in the deflected angles of geodesics terminating at infinity, as observed from different inclination angles $\theta_{0}$ and plotted against observation time $t_{0}-r_{*}(r_{o})$. The impact parameter is fixed at $\tilde{\rho}=10M$ on the $+\tilde{X}$ axis.
We also display cases with different observer inclination angles $\theta_{o}$ in Fig. 4, fixing the impact parameters at $\tilde{\rho}=10M$ along the $+\tilde{X}$-axis. Similar to the observations in Fig. 3, the variations exhibit two types of behavior: a slow decay predominates at smaller $\theta_{o}$, while larger $\theta_{o}$ are primarily characterized by ringdown-like decay and oscillation. This distinction can again be attributed to the propagation of the polar angle from infinity to the BH, where smaller $\theta_{o}$ corresponds to larger $\theta$ in this trajectory. For $\theta_{o}=\pi(1-10^{-4})$ for example, the deflection angle mimics the ringdown frequency to one part in $10^{10}$!
@@
Furthermore, we consider the effect of incorporating the QNM’s mirror mode with equal energy, reflecting a more realistic scenario where the SMBHB was in a quasi-circular orbit before the merger. This setup also exhibits the two previously discussed characteristics: an initial rapid decay during the ringdown phase, followed by a late-time power-law tail. Notably, the ringdown phase becomes more pronounced as the metric perturbations are now comparable at both $\theta\sim 0$ and $\theta\sim\pi$.
@@ IV Discussion
This study has explored the dynamic lensing of photons during the ringdown phase following the merger of SMBHBs. By calculating photon trajectories around a relaxing Kerr BH using backward ray tracing, we demonstrate how metric perturbations during the ringdown significantly influence photon paths. In some instances, variations in photon deflection angles closely correspond to the GW ringdown behavior, particularly when photons co-propagate with GWs from the BH to the observer, thereby offering insights into the fascinating dynamics of spacetime.
@@
The visualization of ringdown through photon lensing opens a new avenue for multi-messenger astronomy. While pulsar timing arrays Agazie et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023), astrometry Braginsky et al. (1990); Book and Flanagan (2011); Moore et al. (2017), binary orbits Blas and Jenkins (2022a, b), and fast radio burst timing Lu et al. (2024) may directly detect ringdown GWs from SMBHB mergers, observations from the Event Horizon Telescope Akiyama et al. (2019, 2022) and future upgrades Ayzenberg et al. (2023), including space missions Haworth et al. (2019); Johnson et al. (2024); Akiyama et al. (2024), could provide visualizations of a ‘Waltz’ between GW and photon co-propagation. Each pixel on the observer plane could reveal additional insights by dissecting the metric perturbations in the vicinity of the relaxing BH.
Clearly, even though we focus on ringdown, their quasi-periodic nature indicates that our results can be taken as indicative of what happens during earlier stages in the life of a binary emitting gravitational radiation.
While our use of backward ray tracing sometimes leads to the disappearance or replacement of a new source at a given pixel, predicting the apparent position shift of a source remains challenging. A future research direction will involve studying photon propagation from a given source point towards the observer in a dynamic spacetime. This endeavor will require a novel approach termed forward ray tracing, presently only feasible by leveraging conserved quantities in Kerr spacetime Zhou et al. (2024).
@@ Acknowledgements.
We thank Maarten van de Meent for helpful discussions and for providing us with useful routines for metric reconstruction. We also extend our thanks to Michael Boyle for his insightful discussions on spin-weighted spherical harmonics. Z.Z. acknowledges financial support from China Scholarship Council (No. 202106040037). V.C. is a Villum Investigator and a DNRF Chair. We acknowledge support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. We acknowledge financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101007855 and No 101131233. Y.C. acknowledges the support of the Rosenfeld foundation and the European Consortium for Astroparticle Theory in the form of an Exchange Travel Grant.
@@ Appendix A Metric Perturbations During Ringdown
A.1 Newman-Penrose formalism
For completeness, we review the Newman-Penrose (NP) formalism in this section, adopting the convention from Refs. van de Meent (2016); Spiers et al. (2023); Pound and Wardell (2020). Note the overall sign difference in some NP quantities compared to the common convention, as seen, for example, in Ref. Teukolsky (1973). This ensures that the background values remain consistent with cases where the metric signature is (+,,,(+,-,-,-)( + , - , - , - ). In the NP formalism, four null vectors $e_{(\mu)}^{a}$ are defined as
$e_{(\mu)}^{a}=\left\{e_{(0)}^{a},e_{(1)}^{a},e_{(2)}^{a},e_{(3)}^{a}\right\}\equiv\left\{l^{a},n^{a},m^{a},\bar{m}^{a}\right\}$ (1)
@@ A.1 Newman-Penrose formalism
$l^{a}n_{a}=-1\quad\text{and}\quad m^{a}\bar{m}_{a}=1\,.$ (2)
@@
$\hat{D}\equiv l^{a}\nabla_{a},\,\hat{\Delta}\equiv n^{a}\nabla_{a},\,\hat{\delta}\equiv m^{a}\nabla_{a},\,\hat{\bar{\delta}}\equiv\bar{m}^{a}\nabla_{a}.$ (4)
@@
$\gamma_{(a)(b)(c)}\equiv g_{\mu\lambda}e_{(a)}^{\mu}e_{(c)}^{\nu}\nabla_{\nu}e_{(b)}^{\lambda}\,,$ (5)
@@
$\kappa$ $\equiv-\gamma_{(3)(1)(1)}\,,\,\varpi\equiv-\gamma_{(2)(4)(1)}\,,\,\epsilon\equiv-\frac{1}{2}\left(\gamma_{(2)(1)(1)}+\gamma_{(3)(4)(1)}\right),$ (6)
@@
$\mathcal{L}=e_{(\mu)}^{a}\nabla_{a}\,$ (7)
can be written as Wald (1978); Dias et al. (2009); Cartas-Fuentevilla (1997)
$\mathcal{L}^{\dagger}=-\mathcal{L}-\nabla_{a}e_{(\mu)}^{a}$ (8)
@@
$\hat{D}^{\dagger}$ $=-(\hat{D}+\varepsilon+\bar{\varepsilon}-\rho-\bar{\rho})\,,\,\hat{\Delta}^{\dagger}=-(\hat{\Delta}-\gamma-\bar{\gamma}+\mu+\bar{\mu}),$ (9)
@@
$e^{\mu}_{\mathrm{(1)}}=l^{\mu}$ $=\frac{1}{\Delta_{\mathrm{BL}}}\left(r^{2}+a^{2},\Delta_{\mathrm{BL}},0,a\right)\,,\quad e^{\mu}_{\mathrm{(2)}}=n^{\mu}=\frac{1}{2\Sigma_{\mathrm{BL}}}\left(r^{2}+a^{2},-\Delta_{\mathrm{BL}},0,a\right),$ (10)
@@
$\Delta_{\mathrm{BL}}\equiv r^{2}-2Mr+a^{2}\,,\quad\Sigma_{\mathrm{BL}}\equiv r^{2}+a^{2}\cos^{2}\theta\,.$ (11)
@@
$\kappa=\lambda=\nu=\sigma=\epsilon=0\,,$ (12)
@@
$\rho=-1/(r-ia\cos\theta)\,,\,\beta=-\bar{\rho}\cot\theta/(2\sqrt{2})\,,\,\varpi=ia\rho^{2}\sin\theta/\sqrt{2}\,,$ (13)
@@
The Wely scalars are defined as
$\Psi_{0}\equiv C_{(1)(3)(1)(3)}=C_{\mu\nu\rho\sigma}l^{\mu}m^{\nu}l^{\rho}m^{\sigma}\,,\,\Psi_{1}\equiv C_{(1)(2)(1)(3)}=C_{\mu\nu\rho\sigma}l^{\mu}n^{\nu}l^{\rho}m^{\sigma}\,,$ (14)
@@
$\hat{D}\Psi_{2}=3\rho\Psi_{2}\,,\,\hat{\Delta}\Psi_{2}=-3\mu\Psi_{2}\,,\,\hat{\delta}\Psi_{2}=3\tau\Psi_{2}\,,\,\hat{\bar{\delta}}\Psi_{2}=-3\varpi\Psi_{2}\,.$ (16)
A.2 Metric Reconstruction
Cohen, Chrzanowski, and Kegeles (CCK) formulated a method for reconstructing vacuum metric perturbations within a radiation gauge Chrzanowski (1975); Cohen and Kegeles (1974); Kegeles and Cohen (1979). Here, we briefly review the procedure of CCK reconstruction, which is detailed in Ref. Toomani et al. (2022); Spiers et al. (2023); Green et al. (2020); Campanelli and Lousto (1999); Whiting and Price (2005); Stewart (1979). Many of our calculations heavily depend on the Newman-Penrose (NP) formalism, adhering to the conventions outlined in Appendix A.1.

### v3

@@ III Lensing Tomography
Figure 4: Top: Deflection angle of geodesics terminating at infinity, $\tilde{\delta}$, defined as the grand arc difference from the unperturbed direction, plotted as a function of observation time $\tilde{t}_{o}$. Different colors represent geodesics launched with varying relative impact parameters $\Delta\tilde{\rho}$ along the $+\tilde{X}$-axis. The BH, observer, and ringdown configuration are consistent with the setup described in Fig. 2. Early stages show a dominant exponential decay and oscillation behavior, in very good agreement with the GW ringdown depicted in the bottom panel. The envelope of the $\Delta\tilde{\rho}/M=0.5$ curve is fitted by $\tilde{\delta}\sim A_{\tilde{\delta}}\exp(\mathrm{Im}(\omega)\,\tilde{t}_{o})$, illustrated by the dot-dashed line. At later times, a slow decay appears, fitted by $\tilde{\delta}\propto\tilde{t}_{o}^{-3}$, shown by the dashed line. Bottom: Amplitudes of the GWs in the two polarizations, $h_{+}$ and $h_{\times}$, in transverse traceless gauge, as observed by an observer at $\theta_{o}/\pi=1-10^{-4}$.
Note that the scaling $A_{\tilde{\delta}}\propto 1/\Delta\tilde{\rho}$ differs from the inverse-cubic dependence reported in Refs. [9, 51, 52, 53]. This distinction arises because, in our case, the dominant deflection occurs locally when the photon first encounters the ringdown perturbation near $t\approx r_{*}$ – close to the BH and at early times – when the photon and GW propagation directions are not yet fully aligned. As a result, the deflection, which is proportional to the contraction $h_{\mu\nu}p^{\mu}p^{\nu}$, is primarily governed by the $1/r$ component of the metric perturbation, rather than the $1/r^{3}$ contribution. In contrast, the inverse-cubic scaling in Refs. [9, 51, 52, 53] assumes a source that emits GWs continuously and nearly isotropically long before the photon arrives, such that only the component of the GW field with momentum aligned with the photon direction contributes appreciably to the net deflection.
Furthermore, as the impact parameter approaches the critical curve, specifically when $\Delta\tilde{\rho}\ll M$, there is a noticeable increase in the overall amplitude of $\tilde{\delta}$. This enhancement can be attributed to the instability of the BH light ring region, which causes an exponential growth in the perturbed deviation during propagation around the BH [40, 41, 42, 43, 44, 45, 46, 47, 48, 32, 50].
On the other hand, at late times, the evolution exhibits a power-law decay. This distinction arises from deflections developing in two distinct regions. During the early stages, the photon encounters the GW field close to the BH, where the $1/r$ component of the perturbation dominates and induces a deflection that decays locally in an exponential fashion—reflecting the ringdown behavior of the GW itself. At late times, the encounter occurs farther from the BH, where the photon and GW propagate nearly anti-parallel. In this regime, the deflection effect receives non-vanishing contributions only from the $1/r^{3}$ component of the metric perturbations. Indeed, our results are consistent with a power-law scaling, $\tilde{\delta}\propto\tilde{t}_{o}^{-3}$, as shown by the dashed fit in Fig. 4, following the behavior predicted in Refs. [9, 51, 52, 53].
Figure 5: Evolutions of the deflection angles of geodesics terminating at infinity, observed from various inclination angles $\theta_{0}$ and plotted against observation time $\tilde{t}_{o}$. The impact parameter is consistently set at $\tilde{\rho}/M=10$ along the $+\tilde{X}$ axis. The dashed black line indicates that $\tilde{\delta}$ scales proportionally to $\tilde{t}_{o}^{-3}$. Inset: The relative strain amplitudes of the GWs, $A_{h}$, defined as $h\equiv(h_{+}^{2}+h_{\times}^{2})^{1/2}\sim A_{h}\exp(\mathrm{Im}(\omega)\tilde{t}_{o})$, are plotted as a function of the $\theta_{o}$. The amplitude is most prominent at $\theta_{o}=\pi$.
We also present cases with different observer inclination angles $\theta_{o}$ in Fig. 5, with the impact parameters fixed at $\tilde{\rho}/M=10$ along the $+\tilde{X}$-axis. Similar to the observations in Fig. 4, the evolutions exhibit two distinct types of behavior: a ringdown-like exponential decay with oscillations, and a subsequent power-law decay proportional to $\tilde{t}_{o}^{-3}$. The power-law tail becomes more pronounced at smaller $\theta_{o}$. This effect is due to the polar angle $\theta$ of photon geodesics at $t\approx r_{*}$ being closer to $\pi$, where the metric perturbation of the $\ell=m=2$ mode is most dominant, as shown in the angular distribution of the GW ringdown in the inset panel. For $\theta_{o}=\pi(1-10^{-4})$, the deflection angle is predominantly characterized by ringdown behavior, with the oscillation frequency mirroring the ringdown frequency to one part in $10^{10}$.
@@
Furthermore, we consider the effect of incorporating the QNM’s mirror mode with equal energy, reflecting a more realistic scenario where the SMBHB was in a quasi-circular orbit before the merger. This setup also exhibits the two previously discussed characteristics: an initial rapid decay during the ringdown phase, followed by a late-time power-law tail, as shown in Supplemental Material C [18].
@@ IV Discussion
This study explored the dynamic lensing of photons during the ringdown phase after SMBHB mergers. By calculating photon trajectories around a relaxing Kerr BH using backward ray tracing, we demonstrate how metric perturbations during the ringdown significantly influence photon paths. In most cases, the evolutions of photon deflection angles initially exhibit behavior akin to gravitational wave ringdown, providing insights into the fascinating dynamics of spacetime.
@@
The visualization of ringdown through photon lensing opens a new avenue for multi-messenger astronomy. While pulsar timing arrays (PTAs) [54, 55, 56, 57], astrometry [58, 59, 60], binary orbits [61, 62], and fast radio burst timing [63] may directly detect ringdown GWs from SMBHB mergers, observations from the Event Horizon Telescope (EHT) [64, 65] and future upgrades [66], including space missions [67, 68, 69], could provide visualizations of a ‘Waltz’ between GW and photon co-propagation. Each pixel on the observer plane could reveal additional insights by dissecting the metric perturbations in the vicinity of the relaxing BH.
A critical question is the feasibility of observing ringdown in electromagnetic channels. This requires the telescope’s Gaussian kernel, represented by $\theta_{\rm res}/2.4$ (where $\theta_{\rm res}$ is the angular resolution), to be finer than the changes in the impact parameter on the observer plane, scaling as $\tilde{\delta}\times r_{s}/r_{o}$ with $r_{s}$ being the distance to the BH. Substituting from Eq. (5) and using $E_{\mathrm{tot}}\sim 0.27Mq^{2}/(1+q)^{4}$, where $q$ is the mass ratio [37], we derive the requirement:
$\frac{M}{r_{o}}>0.67\,\theta_{\mathrm{res}}\beta\left(2+q+\frac{1}{q}\right),$ (6)
where $\beta\equiv\Delta\tilde{\rho}/r_{s}<1$ is a geometric factor dependent on the configuration of the light source, the BH, and the observer. The current angular resolution achievable by the EHT is $20\mathrm{\mu as}$, anticipated to improve to $3\mathrm{\mu as}$ with space-based missions like the Black Hole Explorer (BHEX) [68, 70, 71, 67]. Further enhancements in resolution can be achieved by extending the baselines to locations such as the Moon or the second Sun-Earth Lagrange point [48].
For a merger involving BHs with comparable mass ratios, assuming $M=10^{10}\,M_{\odot}$ as suggested by PTA observations [72], BHEX could detect lensing signatures from distant sources out to $r_{o}\sim 12~{}\mathrm{Mpc}/\beta$. For well-monitored SMBHs such as Sgr A and M87, even mergers with smaller mass ratios become observable, enabling detection of events with $q$ as small as $0.5\beta$. A source with a small $\beta$—typically located on the far side of the BH, such as a jet feature, an accretion flow near the light ring, or a background star—can therefore significantly enhance the detection prospects.
In the case of background stars lensed by Sgr A, Ref. [73] predicts that upcoming telescopes like ELT, TMT, and GMT will detect $\mathcal{O}(100)$ stars with angular separations of $\sim$1 mas and distances of $\sim$1 pc behind the BH, corresponding to $\beta\sim 10^{-5}$. Given their mas-level angular resolution, these telescopes could detect lensing signatures from intermediate-mass BHs merging into Sgr A with mass ratios as small as $q\sim 10^{-3}$.
While our focus here is on ringdown, the underlying formalism applies more broadly to dynamic metric perturbations. For instance, a single SMBH surrounded by a bosonic superradiant cloud could generate sustained periodic perturbations, potentially observable by EHT-class instruments as time-dependent lensing signatures in the accretion flow near the photon ring [50]. Moreover, our approach is applicable to the long-lasting inspiral stage of a binary system, where the accumulation of signal-to-noise ratio could make such searches feasible. A more precise prediction of the apparent position shift of a point source would require generalizing forward ray tracing [74] in perturbed spacetime.
@@ Acknowledgements.
We thank Maarten van de Meent for helpful discussions and for providing us with useful routines for metric reconstruction. We also extend our thanks to Michael Boyle for his insightful discussions on spin-weighted spherical harmonics. Z.Z. acknowledges financial support from China Scholarship Council (No. 202106040037). The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. V.C. and Y. C. acknowledge support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. V.C. is a Villum Investigator and a DNRF Chair. V.C. acknowledges financial support provided under the European Union’s H2020 ERC Advanced Grant “Black holes: gravitational engines of discovery” grant agreement no. Gravitas–101052587. Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101007855 and No 101131233. Y.C. is supported by the Rosenfeld foundation in the form of an Exchange Travel Grant and by the COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
@@ Appendix A Metric Perturbations During Ringdown
1 Newman-Penrose formalism
For completeness, we review the Newman-Penrose (NP) formalism in this section, adopting the convention from Refs. [97, 89, 88]. Note the overall sign difference in some NP quantities compared to the common convention, as seen, for example, in Ref. [16]. This ensures that the background values remain consistent with cases where the metric signature is (+,,,(+,-,-,-)( + , - , - , - ). In the NP formalism, four null vectors $e_{(\mu)}^{a}$ are defined as
$e_{(\mu)}^{a}=\left\{e_{(0)}^{a},e_{(1)}^{a},e_{(2)}^{a},e_{(3)}^{a}\right\}\equiv\left\{l^{a},n^{a},m^{a},\bar{m}^{a}\right\}$ (A.1)
@@ 1 Newman-Penrose formalism
$l^{a}n_{a}=-1\quad\text{and}\quad m^{a}\bar{m}_{a}=1\,.$ (A.2)
@@
$\hat{D}\equiv l^{a}\nabla_{a},\,\hat{\Delta}\equiv n^{a}\nabla_{a},\,\hat{\delta}\equiv m^{a}\nabla_{a},\,\hat{\bar{\delta}}\equiv\bar{m}^{a}\nabla_{a}.$ (A.4)
@@
$\gamma_{(a)(b)(c)}\equiv g_{\mu\lambda}e_{(a)}^{\mu}e_{(c)}^{\nu}\nabla_{\nu}e_{(b)}^{\lambda}\,,$ (A.5)
@@
$\kappa$ $\equiv-\gamma_{(3)(1)(1)}\,,\,\varpi\equiv-\gamma_{(2)(4)(1)}\,,\,\epsilon\equiv-\frac{1}{2}\left(\gamma_{(2)(1)(1)}+\gamma_{(3)(4)(1)}\right),$ (A.6)
@@
$\mathcal{L}=e_{(\mu)}^{a}\nabla_{a}\,$ (A.7)
can be written as [93, 81, 78]
$\mathcal{L}^{\dagger}=-\mathcal{L}-\nabla_{a}e_{(\mu)}^{a}$ (A.8)
@@
$\hat{D}^{\dagger}$ $=-(\hat{D}+\varepsilon+\bar{\varepsilon}-\rho-\bar{\rho})\,,\,\hat{\Delta}^{\dagger}=-(\hat{\Delta}-\gamma-\bar{\gamma}+\mu+\bar{\mu}),$ (A.9)
@@
$e^{\mu}_{\mathrm{(1)}}=l^{\mu}$ $=\frac{1}{\Delta_{\mathrm{BL}}}\left(r^{2}+a^{2},\Delta_{\mathrm{BL}},0,a\right)\,,\quad e^{\mu}_{\mathrm{(2)}}=n^{\mu}=\frac{1}{2\Sigma_{\mathrm{BL}}}\left(r^{2}+a^{2},-\Delta_{\mathrm{BL}},0,a\right),$ (A.10)
@@
$\Delta_{\mathrm{BL}}\equiv r^{2}-2Mr+a^{2}\,,\quad\Sigma_{\mathrm{BL}}\equiv r^{2}+a^{2}\cos^{2}\theta\,.$ (A.11)
@@
$\kappa=\lambda=\nu=\sigma=\epsilon=0\,,$ (A.12)
@@
$\rho=-1/(r-ia\cos\theta)\,,\,\beta=-\bar{\rho}\cot\theta/(2\sqrt{2})\,,\,\varpi=ia\rho^{2}\sin\theta/\sqrt{2}\,,$ (A.13)
@@
The Weyl scalars are defined as
$\Psi_{0}\equiv C_{(1)(3)(1)(3)}=C_{\mu\nu\rho\sigma}l^{\mu}m^{\nu}l^{\rho}m^{\sigma}\,,\,\Psi_{1}\equiv C_{(1)(2)(1)(3)}=C_{\mu\nu\rho\sigma}l^{\mu}n^{\nu}l^{\rho}m^{\sigma}\,,$ (A.14)
@@
$\hat{D}\Psi_{2}=3\rho\Psi_{2}\,,\,\hat{\Delta}\Psi_{2}=-3\mu\Psi_{2}\,,\,\hat{\delta}\Psi_{2}=3\tau\Psi_{2}\,,\,\hat{\bar{\delta}}\Psi_{2}=-3\varpi\Psi_{2}\,.$ (A.16)
2 Metric Reconstruction
Cohen, Chrzanowski, and Kegeles (CCK) formulated a method for reconstructing vacuum metric perturbations within a radiation gauge [26, 27, 28]. Here, we briefly review the procedure of CCK reconstruction, which is detailed in Ref. [92, 89, 82, 77, 94, 90]. Many of our calculations heavily depend on the Newman-Penrose (NP) formalism, adhering to the conventions outlined in Appendix 1.